A plane travels 1050 miles with a tail wind in 2 hours and 30 minutes and returns in 3 houts and 30 minutes against the head wind. Find the speed of the plane in still air and the wind velocity.
Answer by lwsshak3(11628)

A plane travels 1050 miles with a tail wind in 2 hours and 30 minutes and returns in 3 hours and 30 minutes against the head wind. Find the speed of the plane in still air and the wind velocity.
***
let x=speed of plane in stll air
let c=speed of wind
x+c=speed of plane with tail wind
x-c=speed of plane with head wind
travel time=distance/speed
..
1050/(x+c)=2.5
1050/(x-c)=3.5
..
2.5x+2.5c=1050
3.5x-3.5c=1050
..
8.75x+8.75c=3675 (mult eq. by 3.5)
8.75x-8.75c=2625 (mult eq. by 2.5)
add
17.5x=6300
x=360
2.5c=1050-2.5x=1050-900=150
c=60
speed of plane in stll air=360 mph
speed of wind=60 mph